OpenStudy (anonymous):

someone help? what is f(-3) if f(x)=|2x-1|-5

OpenStudy (texaschic101):

f(-3) = | 2(-3) - 1 | - 5 f(-3) = | -6 - 1| - 5 f(-3) = | -7 | - 5 --- absolute values are positive even when they say negative f(-3) = 7 - 5 f(-3) = 2

OpenStudy (anonymous):

ah thank you again so so much help❤️